A population consist of the numbers 3,5,10,11 and 6. Let us list all possible samples of size 3 fr this population and compute the mean of each sample

1
Expert's answer
2021-11-29T19:21:44-0500

We have given the population "3,5,10,11\ and \ 6" and "5."

Sampling distribution of size 3 can be given as






Sampling distribution of sample mean,

"P(\bar X=4.66) = \dfrac{1}{5} ," "P(\bar X=3.66) = \dfrac{1}{10},"

"P(\bar X=6.66) = \dfrac{1}{10}," "P(\bar X = 7.66) = \dfrac{1}{10},"

"P(\bar X=5.66)= \dfrac{1}{10}"

"P(\bar X=5) = \dfrac{1}{5}"

"P(\bar X=7.33) = \dfrac{1}{10}"

"P(\bar X=6.33) = \dfrac{1}{10}"